OpenStudy (anonymous):

find the required number TU=3 RS=15 TW=4 TR=9 UW=

OpenStudy (cherylim23):

UW = UT + TW = -3 + 4 = 1?

OpenStudy (anonymous):

here is a photo to help explain better

OpenStudy (cherylim23):

I thought it was vectors, well... 3/9 = UW/15 Solve that and UW = 5

OpenStudy (anonymous):

\[\frac{TU}{TR}=\frac{UW}{RS}\] \[\frac{3}{9}=\frac{UW}{15}\] Cross multiply: 45=9(UW)

OpenStudy (anonymous):

what about This one. What is WS